OpenStudy (anonymous):

The work of a student trying to solve the equation 2(2x – 3) = 6 + 2x + 4 is shown below: Step 1: 2(2x – 3) = 6 + 2x + 4 Step 2: 4x – 3 = 10 + 2x Step 3: 4x – 2x = 10 + 3 Step 4: 2x = 13 Step 5: x = 6.5 In which step did the student first make an error and what is the correct step? Step 2: 4x – 6 = 10 + 2x Step 2: 4x – 3 = 2 + 2x Step 3: 4x + 2x = 10 + 3 Step 3: 4x – 2x = 10 - 3

sammixboo (sammixboo):

What do you think? Remember first step is to always simplify. Did they simplify right?

OpenStudy (anonymous):

i am just making shure my questions are right

OpenStudy (anonymous):

It's Step 2.

OpenStudy (gorv):

first check step 1 ....is it correct ???

OpenStudy (gorv):

@agl202304

OpenStudy (anonymous):

im pretty shure but i don't know!!

OpenStudy (gorv):

yeah it is correct .....step 1 is same as equation in our Q

OpenStudy (gorv):

now step 2 while we open the bracket we multiply the term outside with each term like 2*(3+5)=2*3+5*3

OpenStudy (gorv):

can u tell that he did step 2 right or wrong??

OpenStudy (anonymous):

there is two step2s and two step 3s wich one is it?

OpenStudy (gorv):

open the bracket 2*(2x-3)=2*2x-2*3 2*2x-2*3=10+2x tell me which one is correct??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

i need to submit my assingment so can you plz tell me?

OpenStudy (gorv):

multiply 2 4x-6=10+2x

OpenStudy (anonymous):

i got 24

OpenStudy (gorv):

how u will get 24 .??

OpenStudy (anonymous):

plz give me the answer?

OpenStudy (gorv):

lol i already explained 4x-6=10+2x